a particel executing shm has an amplitude of 6cm its acceleration at a distance 2cm from mean position is 8cm .the maximum speed is?

Grade:12

1 Answers

Ashutosh Mohan Sharma
7 years ago
a= (ω^2)*y
(ω^2)*2= 8
(ω^2) =4 so ω=2rads per second
maxV=A*ω= 6*2=12 cm per secpnd